OpenStudy (anonymous):

Two perpendicular lines intersect at the origin. If the slope of the first line is , what is the equation of the second line? A y=-2x b y=1/2x c y=-1/2x d y=2x

OpenStudy (anonymous):

the first slope is -1/2

OpenStudy (souvik):

for perpendicular lines..m1*m2=-1 here m1=-1/2 so m2=2 so the equation of the line is...(y-y1)=m2*(x-x1) the line goes through (0,0) so y1=0 and x1=0 the equation is y=2x